SecureProtectContain:

Simplify. Work is expected as this is the written part of the course. Then, submit through your student upload folders. df cos4theta - 2cos2thetha + 1 sint+cottcost Solve for theta. 12=9-6costheta

justjm:

1. \(\cos^{4}\theta-2\cos^{2}\theta+1\) Let \(\text{x}=cos^2\theta \) \(x^{2}-2x+1\) You will find this is a perfect square trinomial. Simplify \((x-1)^2\) Plug x back in \(\left(\cos^{2}\theta-1\right)^{2}\) Use the identity that \(sin^2 u + cos^2 u = 1\) This is rearranged to \(cos^2u-1 = -sin^2u\) \((-sin^2\theta)^2\) Simplify again \(sin^4\theta\) 2. \(sin(t)+cot(t)cos(t)\) \(sin(t)+\frac{cos(t)}{sin(t)}cos(t)\) \(\sin\left(t\right)+\frac{\cos^{2}\left(t\right)}{\sin\left(t\right)}\) Find their common denominators \(\frac{\sin^{2}\left(t\right)+\cos^{2}\left(t\right)}{\sin\left(t\right)}\) Use the identity that \(sin^2 u + cos^2 u = 1\) \(\frac{1}{sin(t)}\) And as you know with the reciprocal identity, \(\frac{1}{sin(u)}=csc(u)\) Then you can do that. 3. You would do basic algebra at first \(12=9-6\cos\theta\) \(3=-6\cos\theta\) \(-\frac{1}{2}=\cos\left(\theta\right)\) Take cos inverse \(\cos^{-1}\left(-\frac{1}{2}\right)=\theta\) If you know your unit circle well, cos is -1/2 at 2pi/3. But since your question didn't tell any bounds, we can put the answer as \(\frac{2\pi}{3}+2n\pi~\) where n is an integer

justjm:

By the way I believe your username is a bit wrong lol. I think you meant it to be Secure *Contain Protect* if you want it to mean the SCP foundation xD

SecureProtectContain:

oh I didnt realize I typed it wrong lmao

justjm:

Oh well, I suppose it's SPC foundation now ¯\_(ツ)_/¯

SecureProtectContain:

lol hahaha